Solve 9x (x + 1) = 4
Answer by ReadingBoosters(3246)

9x(x+1)=4
9x^2 + 9x - 4 = 0
(3x + 4)(3x - 1) = 0
...
x = -4/3
x = 1/3
.....................
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.